# Reconstructing jet anisotropies with cumulants

### v1

@@ II.4 Unfolding
To demonstrate the robustness of this procedure, we unfold each $\langle(2k)^{\prime}_{n}\rangle^{\text{reco}}$ corresponding to a model with a given functional form with response matrices constructed from models with different functional forms. This includes samples with with realistic momentum independent $v^{\prime}_{n,\text{jet}}$ functional forms, as well as physically unrealistic $p_{T,jet}$ dependencies on $v^{\prime}_{n,\text{jet}}$ such as $v^{\prime}_{n,\text{jet}}\propto\cos(p_{T,jet})$. The results of this exercise are presented in fig. 1. We show that the corrected $v^{\prime,\text{corr}}_{n,\text{jet}}\{2k\}(p_{T,jet})$ recovers the functional form of a given model, regardless of the functional form present in the construction of the response matrix.
Figure 1: Comparison of the $p_{T}$ dependence of the $v^{\prime}_{2}\{2k\}(p_{T})$ for Au+Au collisions at 200 GeV with several response matrices constructed with different azimuthal distributions of jets. The dashed line represents the true $v^{\prime}_{2}(p_{T})$ values. Response Matrix A, B, and C are constructed with flat constant azimuthal distributions of jets, corresponding to 0.05, 0.1, and 0.2 respectively. Response Matrix D is constructed with a azimuthal distribution that increases linearly with jet $p_{T}$ of jets.
@@ III Results
The simulated data using PYTHIA+TennGen are analyzed using the two- and four-particle cumulant methods with several input flow signals. We present the $v^{\prime}_{n,\text{jet}}\{2\}$ for several $p_{T,jet}$ selection for Au+Au at 200 GeV in Fig. 2. The input $v^{\prime,\text{MC}}_{n,\text{jet}}$ is a constant equal to $0.08$, $0.02$, and $0.01$ for $n=$ 2, 3 and 4, respectively. The $v^{\prime,\text{corr}}_{n,\text{jet}}\{2k\}$ are obtained after five iterations of Bayesian unfolding. Our results reflect the capabilities of using the two-particle cumulant method to reflect the input et $v^{\prime}_{n}$. The presented results also reflected the input sensitivity to the flow harmonic order attenuation.
Figure 2: Comparison of the $p_{T}$ dependence of $v^{\prime}_{2,\text{jet}}(p_{T,jet})$, $v^{\prime}_{3,\text{jet}}(p_{T,jet})$, and $v^{\prime}_{4,\text{jet}}(p_{T,jet})$ for Au+Au collisions at 200 GeV. The dashed line represents the true $v^{\prime}_{n,\text{jet}}(p_{T,jet})$ values.
Extending our analysis to the four-particle cumulant will offer deeper insights into the impact of flow fluctuations on jet flow measurements, as well as the effects of short- and long-range correlations. In this context, short-range correlations primarily arise from localized effects, such as particle decays or intra-jet correlations Lacey (2006); Borghini et al. (2000); Luzum and Ollitrault (2011); Retinskaya et al. (2012); Aad et al. (2012). These contributions can be mitigated by applying multi-particle sub-event cumulant methods. Figure 3 presents the corrected $v^{\prime}_{2,\text{jet}}\{2k\}$ (where $k=$ 1 or 2 for the 2- and 4-particle correlation, respectively) as a function of $p_{T,jet}$ for various functional forms of $v^{\prime,\text{MC}}_{n,\text{jet}}$. These forms encompass different realistic scenarios where $v^{\prime}_{n,\text{jet}}$ remains either independent of or gradually increases with $p_{T,jet}$, as well as physically unrealistic forms, such as $v^{\prime}_{n,\text{jet}}\propto\cos(p_{T,jet})$. The inclusion of both realistic and unrealistic models underscores the robustness of our method in reconstructing differential jet azimuthal anisotropies, even under extreme conditions. Furthermore, our results indicate an agreement between $v^{\prime}_{2,\text{jet}}\{2\}$ and $v^{\prime}_{2,\text{jet}}\{4\}$, reflecting the model’s nature, which lacks flow fluctuations effects.
Figure 3: Comparison of the $p_{T}$ dependence of the $v^{\prime}_{2,\text{jet}}\{2k\}(p_{T})$ for Au+Au collisions at 200 GeV. The dashed line represents the true $v^{\prime}_{n,\text{jet}}(p_{T})$ values.
@@ IV Conclusions
This work demonstrates the utility of two- and four-particle cumulant methods for measuring jet anisotropies, leveraging TennGen and PYTHIA models to simulate data representative jet backgrounds in the quark-gluon plasma. Our findings show that cumulants can be used to effectively reconstruct jet azimuthal anisotropies, capturing their flow characteristics and providing a robust means to analyze high-$p_{T}$ jet azimuthal anisotropy. This methodology offers promise for disentangling long and short range correlations between jets. Since it is not possible to reconstruct an event plane in small systems, this technique can be used to advance the understanding of jet interactions and anisotropy in heavy-ion and smaller collision systems.
Acknowledgements.
@@
Appendix A Cumulant method
We use the method presented in Bilandzic et al. (2011) which calculates multi-particle cumulants in terms of moments of of $\mathbf{Q}$ vectors defined as
$\mathbf{Q}_{n}\equiv\sum_{k=1}^{M}e^{in\phi_{k}}$ (5)
@@ Appendix A Cumulant method
$\langle 2_{n}\rangle=\frac{|\mathbf{Q}_{n}|^{2}-M}{M(M-1)}$ (6)
@@
$\langle 4_{n}\rangle=$ $\frac{|\mathbf{Q}_{n}|^{2}+|\mathbf{Q}_{2n}|^{2}-2\Re\{\mathbf{Q}_{2n}\mathbf{Q}_{n}^{*}\mathbf{Q}_{n}^{*}\}}{M(M-1)(M-2)(M-3)}$ (7)
$-2\frac{(M-2)|\mathbf{Q}_{n}|^{2}-M(M-3)}{M(M-1)(M-2)(M-3)}$
@@
$\langle\langle(2k)_{n}\rangle\rangle=\frac{1}{W\{2k\}}\sum_{i=1}^{N_{\text{events}}}(w\{2k\}){}_{i}(\langle(2k)_{n}\rangle)_{i}$ (8)
where $(w\{2k\})_{i}$ are event weights and $W\{2k\}$ is the sum of all event weights. Common choices for weights are reference multiplicity or particle momentum. In the case when the multiplicity of the reference is limited methods for momentum weighting are outlined in Bilandzic et al. (2011). In our model, $M$ is fixed and the $v_{n}$ of TennGen is independent of multiplicity leading us to use the weights
$w\{2\}=M(M-1)$ (9)
$w\{4\}=M(M-1)(M-2)(M-3)$
@@
The second order cumulant for the reference sub event is defined as
$c_{n}\{2\}\equiv\langle\langle 2_{n}\rangle\rangle,$ (10)
which is simply the event averaged $n^{\mathrm{th}}$-order $2$-particle reference azimuthal correlation defined in Eq. 6. The fourth order reference cumulant is given by
$c_{n}\{4\}\equiv\langle\langle 4_{n}\rangle\rangle-2\cdot\langle\langle 2_{n}\rangle\rangle^{2}.$ (11)
@@
The differential single event azimuthal correlation for jets with respect to the reference correlation is calculated using additional flow vector $\mathbf{p}$, as well as the reference flow vector $\mathbf{Q}$ defined in Eq. 5. To measure the jet $v_{n}$, the Point of Interest (POI) vector $\mathbf{p}$, is built using the azimuthal angle of all jets in an event within $|\eta_{\text{jet}}|<1.1-R$, defined as
$\mathbf{p}_{n}\equiv\sum_{k=1}^{m_{p}}e^{in\psi_{k}}$ (12)
@@
$\langle 2^{\prime}_{n}\rangle=\frac{\mathbf{p}_{n}\mathbf{Q}_{n}^{*}}{m_{p}M}$ (13)
@@
$\langle 4^{\prime}_{n}\rangle$ $=\frac{(|\mathbf{Q}_{n}|^{2}-2M+2)\mathbf{p}_{n}\mathbf{Q}_{n}^{*}-\mathbf{p}_{n}\mathbf{Q}_{n}\mathbf{Q}_{2n}^{*}}{m_{p}M(M-1)(M-2)}$ (14)
where $m_{p}$ is the number of jets within a given $p_{T}$ bin in an event, and $\mathbf{p}_{n}$ is the POI vector. The event averaged $n^{\mathrm{th}}$-order $2k$-particle reference differential azimuthal correlation $\langle\langle(2k)^{\prime}_{n}\rangle\rangle$ is computed similar to Eq. 8 with the corresponding weights
$w^{\prime}\{2\}=m_{p}M$ (15)
$w^{\prime}\{4\}=m_{p}M(M-1)(M-2)$
@@
These definitions for $\langle 2^{\prime}_{n}\rangle$ and $\langle 4^{\prime}_{n}\rangle$ differ slightly from the standard $n^{\mathrm{th}}$-order $2$- and $4$-particle differential single event azimuthal correlations Bilandzic et al. (2011) due to the fact that we do not have to consider points which are labeled as both RFP and POI because the azimuthal angle of the jet axis $\psi_{\text{jet}}$ is a composite angle from recombination of particles using a jet clustering algorithm. In practice, for jet $v_{n}$, it may be necessary to use RFP and POI correlations in different rapidity regions in order to avoid autocorrelations from using jet constituents in the RFP.
@@
$d_{n}\{2\}\equiv\langle\langle 2^{\prime}_{n}\rangle\rangle$ (16)
@@
$d_{n}\{4\}\equiv\langle\langle 4^{\prime}_{n}\rangle\rangle-2\cdot\langle\langle 2^{\prime}_{n}\rangle\rangle\langle\langle 2_{n}\rangle\rangle$ (17)
@@
The reference azimuthal anisotropies are calculated from the reference cumulants $c_{n}\{2k\}$ defined in eqs. 10 11
$v_{n}\{2\}\equiv\sqrt{c_{n}\{2\}}$ (18)
@@
The differential jet azimuthal anisotropies $v^{\prime}_{n,\text{jet}}\{2k\}$ are calculated using the differential cumulants $d_{n}\{2k\}$ defined in eqs. 16 17 with respect to the the reference azimuthal anisotropies $v_{n}\{2k\}$
$v^{\prime}_{n,\text{jet}}\{2\}(p_{T,jet})=\frac{d_{n}\{2\}(p_{T,jet})}{v_{n}\{2\}}$ (20)
$v^{\prime}_{n,\text{jet}}\{2\}(p_{T,jet})=-\frac{d_{n}\{4\}(p_{T,jet})}{(v_{n}\{4\})^{3/4}}$ (21)
where $v^{\prime}_{n,\text{jet}}\{2\}(p_{T,jet})$ and $v^{\prime}_{n,\text{jet}}\{4\}(p_{T,jet})$ are used to denote estimations of the differential flow for jets with a given transverse momentum $p_{T,jet}$ using the $2^{\text{nd}}$ and $4^{\text{th}}$-order differential cumulants respectively.

### v2

@@ II.2 Cumulant Method
To demonstrate the robustness of this procedure, we unfold $\langle(2k)^{\prime}_{n}\rangle^{\text{reco}}$ corresponding to a $p_{T}$ independent $v^{\text{jet}}_{n}$ with response matrices constructed from models with different functional forms. The results of this exercise are presented in Figure 2. We show that the corrected $v^{,\text{jet},\text{corr}}_{n}\{2k\}(p_{T})$ recovers the functional form of a given model, regardless of the functional form present in the construction of the response matrix.
Figure 2: Comparisons of the $v^{\text{jet}}_{2}\{2\}$ calculated with unfolded $\langle 2^{\prime}_{2}\rangle^{\text{corr}}$ for Pythia-8 +TennGen events with a $p_{T}$ independent truth $v_{2}^{\text{jet}}=4\%$, unfolded with response matrices constructed from Pythia-8 +TennGen events where $v_{2}^{\text{jet}}=4\%$ (Response Matrix A), a scaled constant $v_{2}^{\text{jet}}=5\times 4\%$ (Response Matrix B), and a linear $p_{T}$ dependence $v_{2}^{\text{jet}}(p_{T})=4\%+(0.1\%)\times p_{T}$ (Response Matrix C). The dashed line represents the true constant $v^{\text{jet}}_{2}=4\%$ of the unfolded sample.
@@ III Results
The simulated data using Pythia-8 +TennGen are analyzed using the $2$- and $4$-particle cumulant methods with several input flow signals. We present the $v^{\text{jet}}_{n}\{2\}$ for several $p_{T,jet}$ selection for Au+Au at 200 GeV in Figure 3. The input $v^{\text{jet},\text{MC}}_{n,\text{jet}}$ is a constant equal to $0.08$, $0.02$, and $0.01$ for $n=$ 2, 3 and 4, respectively. The $v^{\text{jet},\text{corr}}_{n}\{2k\}$ are obtained after five iterations of Bayesian unfolding. Our results reflect the capabilities of using the $2$-particle cumulant method to reflect the input jet $v^{\text{jet}}_{n}$. The presented results also reflected the input sensitivity to the flow harmonic order attenuation.
Figure 3: The $p_{T}$ dependence of $v^{\text{jet}}_{n}\{2k\}$ for harmonic order $n=2,3$, and $4$, represented by the red squares, purple triangles, and blue circles, respectively. Values calculated using 22- and $4$-particle correlations are differentiated by full ($k=1$) or open ($k=2$) symbols. The dashed line represents the true $v^{\text{jet}}_{n}(p_{T})$ values.
Extending our analysis to the $4$-particle cumulant will offer deeper insights into the impact of flow fluctuations on jet flow measurements, as well as the effects of short- and long-range correlations. In this context, short-range correlations primarily arise from localized effects, such as particle decays or intra-jet correlations [51, 31, 53, 61, 1, 75]. These contributions can be mitigated by applying multi-particle sub-event cumulant methods.
Figure 4 presents the corrected $v^{\text{jet}}_{2}\{2k\}$ (where $k=$ 1 or 2 for the 2- and 4-particle correlation, respectively) as a function of $p_{T,jet}$ for various functional forms of $v^{\text{jet},\text{MC}}_{n}$. These forms encompass different realistic scenarios where $v^{\text{jet}}_{n}$ remains either independent of or gradually increases with $p_{T,jet}$, as well as physically unrealistic forms, such as $v^{\text{jet}}_{n}\propto\cos(p_{T,jet})$. The inclusion of both realistic and unrealistic models underscores the robustness of our method in reconstructing differential jet azimuthal anisotropies, even under extreme conditions. Furthermore, our results indicate an agreement between $v^{\text{jet}}_{2}\{2\}$ and $v^{\text{jet}}_{2}\{4\}$, reflecting the model’s nature, which lacks flow fluctuations effects.
Figure 4: Comparison of the $p_{T}$ dependence of the $v^{\text{jet}}_{2}\{2k\}$ for Au+Au collisions at 200 GeV calculated with 22-particle ($k=1$) and $4$-particle ($k=2$) correlations. The dashed line represents the true $v^{\text{jet}}_{n}(p_{T})$ values. The unfolded $v^{\text{jet}}_{2}\{2k\}$ $v_{2}^{\text{jet}}=4\%$ (a), a linear $p_{T}$ dependence $v_{2}^{\text{jet}}(p_{T})=4\%+(0.1\%)\times p_{T}$ (b), and a sinusoidal $p_{T}$ dependence $v^{\text{jet}}_{n}=4\%\times[2+\cos{0.2p_{T}}]$ (c).
@@ IV Conclusions
This work demonstrates the utility of $2$- and $4$-particle cumulant methods for measuring jet anisotropies, leveraging TennGen and PYTHIA models to simulate data representative jet backgrounds in the quark-gluon plasma. Our findings demonstrate that cumulants can be effectively used to reconstruct jet azimuthal anisotropies, capturing their flow characteristics and providing a robust means to analyze high-$p_{T}$ jet azimuthal anisotropies. This methodology offers promise for disentangling long and short-range correlations between jets. Since it is not possible to reconstruct an event plane in small systems this technique can be used to advance the understanding of jet interactions and anisotropies in heavy-ion and smaller collision systems.
In real heavy-ion collisions, there are several effects, such as uncorrelated jets and resonance decays. These effects could be significant in small systems, and would have to be considered in order to properly interpret the results. Additionally, event-by-event flow fluctuations will contribute to differences in different order cumulants. In A+A systems, where the relative fluctuations are mostly small, $v_{2}\{2\}=\sqrt{v_{2}^{2}+\sigma^{2}}$ and $v_{2}\{4\}\approx v_{2}\{6\}\approx v_{2}\{8\}\approx\sqrt{v_{2}^{2}-\sigma^{2}}$. However, in small systems, the relative fluctuations can be large, and the nature of the underlying distribution becomes important. Addressing these effects will be important for ensuring the applicability of this method to experimental data.
V Acknowledgments
@@
Appendix A Additional Cumulant Method Details
We use the method presented in [29] which calculates multi-particle cumulants in terms of moments of $\mathbf{Q}$ vectors defined as
$\mathbf{Q}_{n}\equiv\sum_{k=1}^{M}e^{in\phi_{k}},$ (8)
@@ Appendix A Additional Cumulant Method Details
$\langle 2_{n}\rangle=\frac{|\mathbf{Q}_{n}|^{2}-M}{M(M-1)},$ (9)
@@
$aligned\langle 4_{n}\rangle=&\frac{|\mathbf{Q}_{n}|^{2}+|\mathbf{Q}_{2n}|^{2}-2\Re\{\mathbf{Q}_{2n}\mathbf{Q}_{n}^{*}\mathbf{Q}_{n}^{*}\}}{M(M-1)(M-2)(M-3)}\\ &-2\frac{(M-2)|\mathbf{Q}_{n}|^{2}-M(M-3)}{M(M-1)(M-2)(M-3)},$ (10)
@@
[…]
where $(w\{2k\})_{i}$ are event weights and $W\{2k\}$ is the sum of all event weights. Common choices for weights are reference multiplicity or particle momentum. In the case when the multiplicity of the reference is limited, methods for momentum weighting are outlined in [29]. In our model, $M$ is fixed, and the $v_{n}$ of TennGen is independent of multiplicity. Therefore we use the weights
$aligned&w\{2\}=M(M-1),\\ &w\{4\}=M(M-1)(M-2)(M-3),\\$ (12)
@@
The second-order cumulant for the reference sub-event is defined as
$aligned c_{n}\{2\}\equiv\langle\langle 2_{n}\rangle\rangle,$ (13)
which is simply the event averaged $n^{\mathrm{th}}$-order $2$-particle reference azimuthal correlation defined in Equation 9. The fourth-order reference cumulant is given by
$aligned c_{n}\{4\}\equiv\langle\langle 4_{n}\rangle\rangle-2\cdot\langle\langle 2_{n}\rangle\rangle^{2},$ (14)
@@
The differential single event azimuthal correlation for jets with respect to the reference correlation is calculated using additional flow vector $\mathbf{p}$, as well as the reference flow vector $\mathbf{Q}$ defined in Equation 8. To measure the jet $v_{n}$, the Point of Interest (POI) vector $\mathbf{p}$ is built using the azimuthal angle of all jets in an event within $|\eta_{\text{jet}}|<1.1-R$, defined as
$\mathbf{p}_{n}\equiv\sum_{k=1}^{m_{p}}e^{in\psi_{k}},$ (15)
@@
$aligned\langle 2^{\prime}_{n}\rangle=\frac{\mathbf{p}_{n}\mathbf{Q}_{n}^{*}}{m_{p}M},$ (16)
@@
$aligned\langle 4^{\prime}_{n}\rangle&=\frac{(|\mathbf{Q}_{n}|^{2}-2M+2)\mathbf{p}_{n}\mathbf{Q}_{n}^{*}-\mathbf{p}_{n}\mathbf{Q}_{n}\mathbf{Q}_{2n}^{*}}{m_{p}M(M-1)(M-2)},$ (17)
where $m_{p}$ is the number of jets within a given $p_{T}$ bin in an event, and $\mathbf{p}_{n}$ is the POI vector. The event averaged $n^{\mathrm{th}}$-order $2k$-particle reference differential azimuthal correlation $\langle\langle(2k)^{\prime}_{n}\rangle\rangle$ is computed similar to Equation 11 with the corresponding weights
$aligned&w^{\prime}\{2\}=m_{p}M,\\ &w^{\prime}\{4\}=m_{p}M(M-1)(M-2),\\$ (18)
@@
These definitions for $\langle 2^{\prime}_{n}\rangle$ and $\langle 4^{\prime}_{n}\rangle$ differ slightly from the standard $n^{\mathrm{th}}$-order $2$- and $4$-particle differential single event azimuthal correlations [29] due to the fact that we do not have to consider points which are labeled as both RFP and POI because the azimuthal angle of the jet axis $\psi_{\text{jet}}$ is a composite angle from recombination of particles using a jet clustering algorithm. In practice, for jet $v_{n}$, it may be necessary to use RFP and POI correlations in different rapidity regions in order to avoid autocorrelations from using jet constituents in the RFP.
@@
$aligned d_{n}\{2\}\equiv\langle\langle 2^{\prime}_{n}\rangle\rangle,$ (19)
@@
$aligned d_{n}\{4\}\equiv\langle\langle 4^{\prime}_{n}\rangle\rangle-2\cdot\langle\langle 2^{\prime}_{n}\rangle\rangle\langle\langle 2_{n}\rangle\rangle,$ (20)
@@
The reference azimuthal anisotropies are calculated from the reference cumulants $c_{n}\{2k\}$ defined in Equations 13 14
$aligned v_{n}\{2\}\equiv\sqrt{c_{n}\{2\}},$ (21)
@@
The differential jet azimuthal anisotropies $v^{\text{jet}}_{n}\{2k\}$ are calculated using the differential cumulants $d_{n}\{2k\}$ defined in Equations 19 20 with respect to the the reference azimuthal anisotropies $v_{n}\{2k\}$
$aligned v^{\text{jet}}_{n}\{2\}(p_{T})=\frac{d_{n}\{2\}(p_{T})}{v_{n}\{2\}},$ (23)
$aligned v^{\text{jet}}_{n}\{2\}(p_{T})=-\frac{d_{n}\{4\}(p_{T})}{(v_{n}\{4\})^{3/4}},$ (24)
where $v^{\text{jet}}_{n}\{2\}(p_{T})$ and $v^{\text{jet}}_{n}\{4\}(p_{T})$ are used to denote estimations of the differential flow for jets with a given transverse momentum $p_{T}$ using the $2^{\text{nd}}$ and $4^{\text{th}}$-order differential cumulants respectively.